We have detected a faulty stick of memory in Karoo which is throwing some errors as such we are scheduling some urgent maintenance to get this resolved to prevent it causing any major issues. This maintenance window will be around 10 minutes whilst the server is powered down, the faulty memory swapped and the server powered back on again but we will update the status page during the process.
Apologies for any disruption this maintenance may cause.
Update @ 6:00PM - The server is now powered down and will be back online in a few minutes.
Update @ 6:08PM - The server is now being powered back on and the memory has been swapped.
Update @ 6:10PM - The server is now back online and we are no longer seeing any memory faults.
We are currently seeing a number of reports of memory errors on both cPanel and on user websites. We are investigating the cause of this issue currently and will provide further updates as soon as possible.
Update @ 9:23AM - We are about to process a manual update and reboot to resolve the ongoing issues with Patagonia, this will cause a few minutes of outage but we will advise once the reboot has begun and once the server is back online.
Update @ 9:31AM - The reboot has started, a further update will be provided once the server is back online.
Update @ 9:36AM - The reboot has now completed and the server is back online as such we will be monitoring the server closely for any further issues.
We're currently investigating disruptions to SQL on websites on the Onyx platform
08:01: This issue has since been resolved.
We are experiencing intermittent issues on the Ordos server today.
We believe this to be resolved now, however, this is still being monitored for further issues which may occur.
No incidents reported
We are currently seeing reports of memory errors and access issues with cPanel on Patagonia, we are investigating the issues and will provide updates shortly.
Update @ 9:29AM - We are about to reboot the server to resolve the ongoing issues, this will take a few minutes to complete but we will update the status once this is completed.
Update @ 9:38AM - The server has been rebooted and the issue appears to be resolved however we will be monitoring the server closely for the next few hours to verify the issue is 100% resolved.
No incidents reported
Our developers are investigating an issue with new sites failing to provision. We'll update this status with any updates - Apologies for the inconvenience.
This has now been resolved - Thank you for your patience.